# RF-DETR: Neural Architecture Search for Real-Time Detection Transformers

### v1

@@ Appendix G Impact of Fixed Architecture on RF100-VL
Table 17: RF100-VL Fixed Architecture Evaluation. We evaluate the transfer of architectures optimized for COCO to RF100-VL. Fixed architecture models perform strongly without additional dataset-specific NAS, with the RF-DETR (large) model achieving the best performance among prior real-time models. However, dataset-specific NAS provides significant further gains.
@@
RF-DETR (Ours) N 30.8M 36.3 2.5 57.6 84.9 62.1 30.7 52.2 66.8
RF-DETR w/ Fine-Tuning (Ours) N 30.8M 36.3 2.5 58.7 85.6 63.5 32.4 52.7 67.0
@@
RF-DETR (Ours) S 33.3M 65.5 3.7 60.7 87.0 66.0 35.4 55.4 69.6
RF-DETR w/ Fine-Tuning (Ours) S 33.3M 65.5 3.7 61.0 87.2 66.4 35.3 55.9 69.8
@@
RF-DETR (Ours) M 33.6M 91.0 4.6 61.5 87.7 67.0 36.44 56.5 69.8
RF-DETR w/ Fine-Tuning (Ours) M 33.6M 91.0 4.6 61.9 87.9 67.3 36.4 56.6 70.1
@@
RF-DETR (Ours) L 34.1M 119.1 6.3 62.3 88.2 68.0 36.4 57.3 70.6
RF-DETR (Ours) w/ Fine-Tuning L 34.1M 119.1 6.3 62.6 88.4 68.2 37.0 57.5 70.5
@@
RF-DETR (Ours) XL 35.0M 199.0 9.8 62.7 88.5 68.5 39.3 58.4 70.4
RF-DETR (Ours) w/ Fine-Tuning XL 35.0M 199.0 9.8 63.1 88.6 69.0 39.6 58.5 70.8
@@
Appendix H Discussion on Notable Discovered Architectures
@@
All “tunable” knobs are used when defining the Pareto-optimal model families, validating our choice of search space. This suggests that expanding the search space may additionally improve downstream performance.
@@
The Pareto optimal models tend to use the same patch size, excluding the Max variants which are just chosen as the highest accuracy running at less than 100ms. For example, the optimal patch size for RF-DETR with a DINOv2-S backbone converges to a size of 16, whereas the DINOv2-B backbone converges to a patch size of 20. The optimal patch size for RF-DETR-Seg with a DINOv2-S backbone is 12. All Pareto-optimal model families simultaneously scale the compute in both the encoder and the decoder. Changing patch size, number of windows, and resolution impacts the encoder, and changing number of decoder layers and number of queries impacts the decoder. For RF-DETR-Seg, scaling resolution also impacts the segmentation head. We find that using 2 windows is typically optimal in the encoder and resolution scales within a family as we increase latency. For the decoder, on COCO the detector model keeps queries constant and scales decoder layers only, while for the segmentation model both are scaled simultaneously. This may be because the depth of the segmentation head is tied to the number of decoder layers, and there may be a certain minimum number of layers in the segmentation head that produce viable masks, so even the lowest latency model uses at least that number of layers. To compensate for the additional latency of those decoder layers, the segmentation model trades off the number of queries it uses to get lower latency for smaller model sizes, which effectively reduces the width of the decoder. While the object detector evaluated on COCO prefers a wide and shallow decoder, the segmentation model prefers a thin and deep decoder.
@@
We find that RF-DETR’s performance depends on the number of spatial locations (e.g. resolution divided by patch size) rather than resolution or patch size individually. Scaling resolution while fixing patch size yields similar results to scaling patch size while fixing resolution, since vision transformers are agnostic to absolute input resolution after the patchify-and-project stage. To verify this, we constructed an alternative family with fixed resolution ($640$) and varied patch sizes to preserve the number of spatial locations. Specifically, we evaluate RF-DETR (nano) with a patch size of $27$, RF-DETR (small) with a patch size of $21$, and RF-DETR (medium) with a patch size of $18$, achieving results nearly identical to the Pareto-optimal family. Notably, patch sizes of $27$ and $18$ were unseen during training, demonstrating RF-DETR’s strong generalization to novel patch sizes.
However, we find that this trend does not hold for RF-DETR-Seg. The segmentation head features are always upsampled to run at 1/4 scale of the input image resolution. Therefore, scaling resolution affects both the number of spatial locations and the segmentation head resolution. Specifically, RF-DETR-Seg (nano) uses a resolution of $312$ and a patch size of $12$, yielding a segmentation head resolution of $78$ with $26$ spatial locations; RF-DETR-Seg (small) uses a resolution of $384$ and a patch size of $12$, yielding a segmentation head resolution $96$ with $32$ spatial locations; and RF-DETR-Seg (medium) uses a resolution of $432$ and a patch size of $12$, yielding a segmentation head resolution of $108$ with $36$ spatial locations. In contrast, scaling patch size alone (e.g., a patch size of $16$ at a resolution of $576$) can keep spatial locations fixed while increasing head resolution (to $144$). This is a more nuanced interaction between patch size and resolution than is observed in RF-DETR object detection. RF-DETR(medium) uses $576/16=36$ spatial locations, while RF-DETR-Seg (medium) uses $432/12=36$ spatial locations, but at a lower resolution, showing that the additional tying of the segmentation mask resolution changes the Pareto-optimal resolution even if the optimal number of spatial locations is the same for a given latency range.
Dataset characteristics influence optimal discovered architectures. We find that when running NAS on RF100-VL datasets, the optimal lower latency models tend to use fewer queries than the COCO models of equivalent latency, which always uses 300. This may be because RF100-VL datasets tend to have fewer objects per image than COCO, so fewer queries are required to find all the objects, as each query is able to find a single object.
@@
Most Pareto-optimal RF-DETRmodels perform best with 2 windows, whereas LW-DETR achieves the best performance with 4 windows. We attribute this difference to how each architecture handles class tokens. LW-DETR’s CAEv2 backbone omits the class token, while RF-DETR’s DINOv2 backbone relies on it as a key part of pre-training. To make windowed attention compatible with class tokens, we duplicate the class token for each window. During global attention, window-level class tokens attend to one another, while all other tokens attend to all class tokens. In practice, RF-DETR (nano), RF-DETR (small), and RF-DETR (medium) all use 2 windows, since duplicating class tokens for additional windows reduces runtime efficiency. As a result, unlike LW-DETR, RF-DETR does not benefit from scaling to 4 windows.
To generate our model based on DINOv2-B, we follow the scaling strategy from ViT-S to ViT-B and just double the width of all layers in the model. We find that this is sufficient to gain strong and differentiated performance when we allow NAS to explore the axis of variation we’ve defined. Notably, different from LW-DETR’s larger variants, we don’t use higher resolution feature maps from the backbone.
Appendix I Visualizing Model Predictions
@@
We visualize model predictions from RF-DETR (nano) and compare them with comparable detection and segmentation baselines in Figure 6. We find that RF-DETR (nano) predicts fewer false positives (e..g mistaking sign post for person). Similarly, RF-DETR-Seg. (nano) predicts more precise object boundaries.

### v2

@@ Appendix I Impact of Fixed Architecture on RF100-VL
Table 18: RF100-VL Fixed Architecture Evaluation. We evaluate the transfer of architectures optimized for COCO to RF100-VL. Fixed architecture models perform well without additional dataset-specific NAS, with the RF-DETR (large) model achieving the best performance among prior real-time models. However, dataset-specific NAS provides significant further gains.
@@
RF-DETR (Ours) N 31.2M 34.5 2.5 57.8 85.1 62.5 30.1 52.2 67.2
RF-DETR w/ Fine-Tuning (Ours) N 31.2M 34.5 2.5 58.6 85.7 63.0 31.0 53.2 67.6
@@
RF-DETR (Ours) S 33.5M 62.4 3.7 60.9 87.5 66.1 34.2 55.7 69.6
RF-DETR w/ Fine-Tuning (Ours) S 33.5M 62.4 3.7 61.2 87.7 66.1 34.9 55.6 69.5
@@
RF-DETR (Ours) M 33.6M 91.0 4.6 61.7 88.0 66.9 35.8 56.5 70.0
RF-DETR w/ Fine-Tuning (Ours) M 33.6M 91.0 4.6 62.0 88.1 67.1 36.2 56.4 70.2
@@
RF-DETR (Ours) L 34.1M 119.1 6.2 62.0 88.1 67.3 36.9 57.1 70.2
RF-DETR (Ours) w/ Fine-Tuning L 34.1M 119.1 6.2 62.3 88.2 67.4 37.1 57.2 70.3
@@
RF-DETR (Ours) XL 35.0M 199.0 9.7 62.6 88.5 67.9 39.0 57.8 70.4
RF-DETR (Ours) w/ Fine-Tuning XL 35.0M 199.0 9.7 63.0 88.7 68.2 38.8 58.2 70.6
@@
Appendix J Ablation on Backbone Architecture with RF20-VL
In Table 19, we reproduce our ablation on the impact of backbone architecture on downstream model performance (Table 6) on RF20-VL. All trends from the main paper hold.
Table 19: Ablation on Backbone RF20-VL.
   w/ CAEv2 ViT/S-16-Truncated Backbone 28.3M 83.7 4.3 64.4 92.2 71.2 33.8 56.8 71.9
   w/ DINOv2 ViT/S-14 Backbone 32.3M 78.2 4.6 65.2 92.5 72.8 37.8 58.2 72.6
   w/ SigLIPv2 ViT/B-32 Backbone 105.1M 81.6 4.5 62.2 91.0 67.8 28.9 54.1 70.3
   w/ SAM2 Hiera-S Backbone 44.0M 109.1 11.2 65.2 92.5 72.7 37.8 58.2 72.1
Appendix K Analysis on Buffering
We further analyze the impact of buffering on the relative ordering of inference speed in Table 20. Notably, we find that buffering beyond 200ms does not change latency measurements. However, we acknowledge that adding a 200ms buffer after every forward pass considerably increases overall inference time. Future work should consider alternatives to buffering to address power throttling.
Table 20: Analysis on Buffering. We evaluate models with different amounts of buffering between consecutive forward passes. We find that buffering beyond 200ms does not provide any additional stability to latency measurements.
    Model     mAP     0 ms     200 ms     400 ms     800 ms
    YOLOv8 (M)     47.3     5.5 ms     5.4 ms     5.4 ms     5.4 ms
    YOLOv11 (M)     48.4     5.0 ms     5.0 ms     5.1 ms     5.1 ms
    RT-DETR (R18)     49.0     4.4 ms     4.4 ms     4.4 ms     4.4 ms
    LW-DETR (M)     52.6     4.5 ms     4.3 ms     4.3 ms     4.3 ms
    D-FINE (M)     54.9     5.7 ms     5.4 ms     5.4 ms     5.4 ms
    RF-DETR (M)     54.7     4.7 ms     4.4 ms     4.4 ms     4.4 ms
Appendix L Discussion on Notable Discovered Architectures
@@
Several trends emerge from our weight-sharing NAS. First, we note that all tunable “knobs” are used when defining Pareto-optimal model families, validating our search space. This suggests that expanding the search space may further improve downstream performance.
@@
Across Pareto-optimal models, patch size is consistent within model families. For example, the optimal patch size for RF-DETRs with a DINOv2-S backbone is 16, RF-DETRs with a DINOv2-B backbone is 20, and RF-DETR-Segs with a DINOv2-S backbone is 12. Pareto-optimal models also jointly scale encoder and decoder compute: patch size, number of windows, and resolution impact the encoder, while decoder depth, and number of queries affect the decoder. For RF-DETR-Seg, scaling resolution impacts the segmentation head. We find that using 2 windows in the encoder is typically optimal and resolution scales within a model family as we increase latency. On COCO, RF-DETR scales decoder depth while keeping the number of queries fixed, while RF-DETR-Seg simultaneously scales both axes. This likely reflects a minimum viable segmentation head depth; to offset its latency, RF-DETR-Seg reduces its total number of queries, yielding a thin, deep decoder, in contrast to RF-DETR’s wide, shallow decoder.
@@
Next, we find that RF-DETR’s performance is more correlated with the total number of spatial locations (e.g. resolution divided by patch size) rather than resolution or patch size alone. Scaling resolution with a fixed patch size yields similar results to scaling patch size with a fixed resolution, since vision transformers are agnostic to absolute input resolution after the patchify-and-project operation. To verify this, we constructed an alternative model family with fixed resolution ($640$) and varied patch sizes to preserve the total number of spatial locations. Specifically, we evaluate RF-DETR (nano) with a patch size of $27$, RF-DETR (small) with a patch size of $21$, and RF-DETR (medium) with a patch size of $18$. Surprisingly, all model results are nearly identical to the Pareto-optimal family. Notably, patch sizes of $27$ and $18$ were unseen during training, demonstrating RF-DETR’s strong generalization to novel patch sizes (Beyer et al., 2023). However, we find that this trend does not hold for RF-DETR-Seg because segmentation features are always upsampled to $\frac{1}{4}$ of the input image resolution. As a result, scaling RF-DETR-Seg’s input resolution affects both the number of spatial locations and the segmentation feature resolution. Specifically, RF-DETR-Seg (nano, small, medium) uses input resolutions of 312, 384, and 432 with patch size 12, yielding segmentation feature resolutions of 78, 96, and 108 and 26, 32, and 36 spatial locations, respectively. In contrast, increasing patch size alone (e.g., patch size 16 at input resolution 576) preserves spatial locations while increasing segmentation feature resolution. As a result, although RF-DETR (medium) and RF-DETR-Seg (medium) both use 36 spatial locations, RF-DETR-Seg operates at lower input resolution, demonstrating that coupling segmentation feature resolution with input resolution shifts the Pareto-optimal operating point.
@@
Further, we find that most Pareto-optimal RF-DETR models perform best with 2 windows, whereas LW-DETR achieves the best performance with 4 windows. We attribute this difference to how each architecture handles class tokens. LW-DETR’s CAEv2 backbone omits the class token, while RF-DETR’s DINOv2 backbone relies on it as a key part of pre-training. To make windowed attention compatible with class tokens, we duplicate the class token for each window. During global attention, window-level class tokens attend to one another, while all other tokens attend to all class tokens. In practice, RF-DETR (nano), RF-DETR (small), and RF-DETR (medium) all use 2 windows, since duplicating class tokens for additional windows reduces runtime efficiency. As a result, unlike LW-DETR, RF-DETR does not benefit from scaling to 4 windows.
Lastly, we note that dataset characteristics influence optimal discovered architectures. We find that the optimal low latency models on RF100-VL tend to use fewer queries than the COCO models of equivalent latency. We attribute this to RF100-VL datasets having fewer objects per image than COCO.
Appendix M Visualizing Model Predictions
@@
We visualize model predictions from RF-DETR (nano) and relevant detection and segmentation baselines in Figure 8. We find that RF-DETR (nano) predicts fewer false positives (e.g. mistaking sign post for person). Similarly, RF-DETR-Seg. (nano) predicts more precise object boundaries.